OpenStudy (anonymous):

How do I factor this: e^(2x)/144 + 9e^(-2x) + 1/2. Is there a method I can use that will work every time?

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@Agent_Sniffles @agent @jiteshmeghwal9 @tkhunny

OpenStudy (anonymous):

well first we know that e^(-2x) is in the form 1/e^(2x) so try to form an expantion\[(x+\frac{1}{x})^2\]

OpenStudy (anonymous):

\[(\frac{e^{2x}}{12}+\frac {3}{e^{2x}})^2\]

OpenStudy (anonymous):

notice that by inspection e^2x is no longer a problem but the numbers 12^2=144 and 3^9

OpenStudy (anonymous):

I'm trying to factor it so I can take the square root of the equation.

OpenStudy (anonymous):

the easy thing about the equation is that we have the form\[(x+\frac{1}{x})^2=x^2+2x \times \frac{1}{x}+\frac{1}{x^2}=x^2+\frac{1}{x^2}+2\] so we let \[e^{2x}=k\] so \[\frac{k^2}{144}+\frac{9}{k^2}+\frac{1}{2}\] and this is \[(\frac{k}{12}+\frac{3}{k})(\frac{k}{12}+\frac{3}{k})=(\frac{k}{12}+\frac{3}{k})^2\] putting back your k=e^2x you get the factor